1. In an open economy, what is the source of supply in the foreign-currency exchange market?

 

A. Net exports

B. Exports

C. Investment and Net Capital outflow

D. Net Capital outflow


Next question:::

 

2. When there is a budget deficit, will net exports increase or decrease?

 

Next question:::

 

3. You have just been hired by the Canadian government to analyze the following scenario. Suppose the Canadian pharmaceutical industry is concerned with foreign prescription drug producers exporting their goods to Canada, a practice that hurts domestic producers. Lobbyists claim that implementing a tariff on imports would shrink the size of the deficit. The following exercise will help you understand how to analyze that claim.

 

A graph is shown with quantity of dollars on the horizontal axis and real exchange rate on the vertical axis. Does demand shift (left/right/stay same) and does the supply curve shift? (left/right/same)

 

Given this change, the value of the dollar (appreciates/depreciates)

 

Now, solve this:

Change to a tariff:
Supply of loanable funds (decrease/no change/increase)

Real interest rate (decrease/no change/increase)

National savings (decrease/no change/increase)

Net exports (decrease/no change/increase)
Answer #1

(1) (D)

In an open economy, net capital outflow is the source of foreign currency supply.

(2) Decrease

By the twin-deficit hypothesis,

T - G = X - M

When (T - G) > 0, There is budget surplus and (X - M) > 0, i.e. Net exports > 0.

When (T - G) < 0, There is budget deficit and (X - M) < 0, i.e. Net exports < 0.

(3)

Demand curve for dollars will shift left (since demand for dollars will fall to pa for imported drugs).

Supply curve for dollars will remain same (since supply of dollars will remain unchanged).

Value of dollar depreciates.

In loanable funds market we use this identity:

S (Savings) = I (Domestic investment) + NCO (Net capital outflow)

Therefore,

Demand for loanable funds = I + NCO

Supply of loanable funds = S

As result of tariff (which increases domestic price and decreases imports, thus increasing net exports and increasing NCO):

Supply of loanable funds - No change

Real interest rate - Increase

National savings - Increase

Net exports - Increase